# Post-selection inference for quantifying uncertainty in changes in variance

### v1

@@ 3.2.2 Estimating $\mathcal{S}$ using Gaussian processes
If we can calculate $\mathcal{S}$ then it is straightforward to calculate the $p$-value. Whilst $\mathcal{S}$ is intractable, as mentioned above, for any value of $\phi$ we can calculate $\mathbb{I}_{\phi\in\mathcal{S}}$. We can represent our beliefs about $\mathbb{I}_{\phi\in\mathcal{S}}$ as a function of $\phi$, through a probability distribution $p(\phi)=\Pr(\mathbb{I}_{\phi\in\mathcal{S}})$ that gives a probability to $\mathbb{I}_{\phi\in\mathcal{S}}$ for each $\phi\in[0,1]$. We can then model this function as a Gaussian process (GP) (see e.g. [37]). This means that we treat $p(\phi)$ as a random function such that for any set of $\phi$ values, $\{\phi_{1},\ldots,\phi_{N}\}$, $\left(p(\phi_{1}),\ldots,p(\phi_{N})\right)$ follows a multivariate Gaussian distribution. We can then take as our estimate, $\Pr(\mathbb{I}_{\phi\in\mathcal{S}})$, the mean of the Gaussian process. Given an observation $\phi\in\mathcal{S}$ or $\phi\notin\mathcal{S}$, the GP will force $p(\phi)$ to take the value either 0 or 1. In between, the mean of the Gaussian process will smooth the function values between these observations.
To implement this, we first need to choose a kernel function, $\boldsymbol{K}(\phi,\phi^{\prime})$, which specifies the covariance between each pair of $\phi$ values. It is natural to choose an exponential kernel $\boldsymbol{K}(\phi_{1},\phi_{2})=e^{-\frac{1}{2l^{2}}|\phi_{1}-\phi_{2}|}$, as this is Markov. (This requires us to choose the value of $l$: we investigate this in Section C.3 of the supplementary material, where we find that at long as $l$ is at least 1, the value we use does not have a noticeable effect on the results.) If we have evaluated $\mathbb{I}_{\phi\in\mathcal{S}}$ at $\phi_{1}<\cdots<\phi_{N}$, and $\phi_{i}<\phi<\phi_{i+1}$ the Markov property means that our Gaussian approximation to $p(\phi)$ will only depend on the values of $\mathbb{I}_{\phi_{i}\in\mathcal{S}}$ and $\mathbb{I}_{\phi_{i+1}\in\mathcal{S}}$. Furthermore, for such a kernel the mean of the Gaussian process will be an average of these values and the prior mean – and thus will always lie within $[0,1]$.
@@
We calculate the estimate of $p(\phi)$ at discrete locations $\boldsymbol{\phi}^{\prime}=\{\phi_{1}^{\prime},\ldots,\phi_{n}^{\prime}\}$, as
$\hat{\boldsymbol{p}}(\boldsymbol{\phi}^{\prime})=\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi}^{\prime})^{\top}\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi})^{-1}\boldsymbol{p}(\boldsymbol{\phi}).$
The most expensive step here computationally is calculating the inverse of $\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi})$, which is an $N\times N$ matrix. However the Markov property of our kernel means that this scales linearly with $N$.
From the output of the Gaussian process, we can estimate the distribution of $\phi$ given $\phi\in\mathcal{S}$ by
@@
$p=\Pr(\phi\leq\phi_{lower}\text{ or }\phi\geq\phi_{upper}~{}|~{}\phi\in\mathcal{S}).$ (11)
@@
Alternatively, we can use importance sampling with $\hat{q}$ as the proposal distribution. This means that we sample $\tilde{\phi}_{1},\ldots,\tilde{\phi}_{\tilde{N}}\sim\hat{q}(\phi)$; then, denoting the true distribution of $\phi$ by $f$, the $p$-value estimate is
$\hat{p}=\frac{\sum_{j=1}^{N}\mathbb{I}\left(\tilde{\phi}_{j}\in\mathcal{S}~{}\&~{}\left(\tilde{\phi}_{j}\leq\phi_{lower}\text{ or }\tilde{\phi}_{j}\geq\phi_{upper}\right)\right)\frac{f(\tilde{\phi}_{j})}{\hat{q}(\tilde{\phi}_{j})}}{\sum_{j=1}^{N}\mathbb{I}\left(\tilde{\phi}_{j}\in\mathcal{S}\right)\frac{f(\tilde{\phi}_{j})}{\hat{q}(\tilde{\phi}_{j})}}.$ (12)
Since we have already calculated $\mathbb{I}_{\phi\in\mathcal{S}}$ for the $\phi$ in our original sample (that we used to fit the Gaussian process), we can also include these points in our importance sampling estimator.
@@ 4 Extensions
In this section we discuss several extensions to the methods we have discussed so far: we consider (i) applying our method to different null hypotheses; (ii) different changepoint detection methods; (iii) reducing the amount of information we condition on using the method of [6].
@@ 4.1 Different null hypotheses
As mentioned in Section 2, we generally want to choose a null hypothesis that there are no changes close to the detected changepoint $\hat{\tau}$. A common way to do this is to use the approach we have used so far, to define $H_{0}$ to be the statement that there is no change within a fixed window $\{\hat{\tau}-h+1,\ldots,\hat{\tau}+h\}$. However, there are modifications or alternatives to this which could be used. For example, we could decide to truncate the region if we find another changepoint (other than $\hat{\tau}$) within the region. Alternatively, if we denote our changepoint of interest $\hat{\tau}_{j}$, we can take the null hypothesis that there is no changepoint within $\{\hat{\tau}_{j-1}+1,\ldots,\hat{\tau}_{j+1}-1\}$ (by convention setting $\hat{\tau}_{0}=0$ and $\hat{\tau}_{K+1}=T$, where $K$ is the number of changepoints). (This approach is considered for the change in mean case, for example, in [20] and [6].)
@@ 4.2 Different changepoint detection methods
We have focused on two approaches to post-selection inference for the change in variance model. The CUSUM approach (Section 3.1) depends on being able to write down a set of linear inequalities in $\phi$ which can be resolved to obtain the set $\mathcal{S}$. We have focused on doing this for binary segmentation, as this is the most straightforward case; however, it is easy to show that we can do the same for variants such as wild binary segmentation, which also involve maximizing $G_{s,e}(t)$ on a series of intervals. In this case, we condition also on the set of random intervals; provided we do this, we can construct linear inequalities in $\phi$ and calculate $\mathcal{S}$ in the same way.
For our second, more general, approach (Section 3.2), we can use any method for changepoint estimation, since we just re-apply the algorithm for each sampled $\phi$. This approach can also be used for the CUSUM case if we use a changepoint algorithm for which we cannot determine $\mathcal{S}$ explicitly.
For some methods it is not necessary to re-run the algorithm from scratch each time. For example, to implement methods such as wild binary segmentation and narrowest over threshold we have to calculate the likelihood ratio statistic on a set of intervals $[s,e]\subseteq[1,T]$. The set of intervals is the same each time we implement the algorithm on $\boldsymbol{X}^{\prime}(\phi)$. Since $\Lambda_{s,e}(t,\phi)$ is constant on intervals which do not overlap with the region of interest (i.e., $e\leq\hat{\tau}-h$ or $s\geq\hat{\tau}+h+1$), for these intervals we only have to calculate $\Lambda_{s,e}(t)$ once. Therefore, at each sampling iteration we only have to re-calculate $\Lambda_{s,e}(t,\phi)$ on intervals which overlap with the region of interest. This can help to speed up the process of estimating $\hat{p}$.
@@ 4.3 Increasing power by conditioning on less information
where here both $\phi$ and $\boldsymbol{W}$ and allowed to vary.
@@ 5.1 Detecting changes using CUSUM
Figure 6: QQ plots of $p$-values obtained using the CUSUM method. We set $T=200$, and sampled data under $H_{0}$ (no change) and with a single changepoint at $\tau=100$. The variance before the change was $\sigma_{1}^{2}=1$ in each case; the variance after the change is $\sigma^{2}\in\{1,0.5,2,4\}$. In each case we estimate a single changepoint using binary segmentation. We calculated $p$-values for different values of $h$, including the case where the window size is equal to the whole data set.
Figure 6 shows QQ plots of $p$-values obtained by sampling from a model with $T=200$. In (a) we sample from $H_{0}$; in (b)–(d) we sample from a model with a single change at $\tau=100$, where in each case the variance before the changepoint is 1, and the variance after the changepoint is 0.5, 2, 4. We see that under $H_{0}$, $p\sim U(0,1)$, and under $H_{1}$ the test has power to detect changes, which increases with the window size $h$ and the size of the change.
@@
Figure 7 shows similar results where we simulate from a model with 4 changes, with $T=200$. In panels (a) and (b) we simulated from a model with 4 equally spaced changes, $\tau=40,80,120,160$. In (c) and (d), changepoints were chosen from a uniform distribution, with the constraint that the minimum distance between changepoints was $h$. In each case we calculated the $p$-value associated with the first changepoint.
@@ 5.2 Detecting changes using the likelihood-ratio test
$N=50$$N=100$$N=200$
Figure 8: QQ plots of $p$-values obtained using the likelihood ratio for three scenarios, using binary segmentation to estimate changepoints. Each line corresponds to the proportion of samples that were used to fit the Gaussian process.
Figures 8 and 9 show QQ plots of estimated $p$-values we obtain on simulated data for three different scenarios: (1) no change; (2) three changes at fixed equidistant locations, $\tau=\frac{T}{4},\frac{T}{2},\frac{3T}{4}$; (3) three changes at random locations, sampled uniformly from $\{h,\ldots,T-h\}$ and constrained to be at least $h$ apart. In each case we set $T=200$ and $h=20$. For the cases with three changes, the variances on each segment were $(1,4,1,0.25)$. For each simulated data set, we estimated changepoints using binary segmentation (Figure 8) or PELT (Figure 9) and calculated an estimate of the $p$-value corresponding to the first detected changepoint using a Gaussian process (as outlined in Section 3.2.2), and using a Gaussian process followed by importance sampling from the posterior distribution, for $N=50,100,200$.
@@
$N=50$$N=100$$N=200$
Figure 9: QQ plots of $p$-values obtained using the likelihood ratio for three scenarios, using PELT to estimate changepoints. Each line corresponds to the proportion of samples that were used to fit the Gaussian process: where this is less than 1, the remaining samples were drawn from the posterior of the Gaussian process.
@@ 6 Application to financial data
Figure 12: Application of our method to financial log returns data. Changepoints are estimated using (i) CUSUM statistic with binary segmentation (top); (ii) likelihood ratio statistic with binary segmentation (middle); (iii) likelihood ratio statistic with PELT (bottom). In each case we estimated 11 changepoints. Vertical lines indicate changepoints: those with a $p$-value smaller than 0.05 (after controlling for multiple testing using the Holm-Bonferroni method) are shown in red, and those with a $p$-value larger than 0.05 in grey.
To investigate how well our method works in a real-world setting, we applied our method to finanical data (Figure 12). The data is log returns from S&P500 stock market data for 2264 days from July 2010 to October 2019111The data was downloaded from https://www.kaggle.com/datasets/mathisjander/s-and-p500-volatility-prediction-time-series-data. We can see from Figure 12 that the data appears to have zero mean and several changes in variance. We estimated changepoints using three different methods: binary segmentation with the CUSUM statistic (top) and the likelihood ratio statistic (middle) and PELT (bottom), and calculated $p$-values for each estimated changepoint. We set thresholds so that in each case we estimated 11 changepoints, and set $h=50$. For the cases where we had to estimate $p$-values, we used a Gaussian process with $N=100$. On each plot estimated changepoints are shown by vertical lines, with red lines corresponding to estimated changepoint locations whose $p$-values were smaller than $0.05$ (after controlling for multiple testing using the Holm-Bonferroni method) and grey lines corresponding to $p$-values which were not significant. Using the likelihood ratio statistic generally seems to give better results than CUSUM, where there are more instances of multiple changepoints being estimated close together. However, these spurious changepoints generally have non-significant $p$-values associated with them, showing that our method for post-selection inference is effective at distinguishing true from false changepoint estimates.
@@ 7 Discussion
The strategies we have developed using Gaussian processes and importance sampling may be useful for other similar situations where we conduct post-selection inference but cannot characterise the selection event explicitly.
Acknowledgement
This work is supported by EPSRC grant number EP/V053590/1.

### v2

@@ 3.2.2 Estimating $\mathcal{S}$ using Gaussian processes
If we can calculate $\mathcal{S}$ then it is straightforward to calculate the $p$-value. For any value of $\phi$ we can represent our belief about $\mathbb{I}_{\phi\in\mathcal{S}}$ by a probability $p(\phi)=\Pr(\mathbb{I}_{\phi\in\mathcal{S}})$. Our approach is to model the function $p(\phi)$ as a Gaussian process (GP) [see e.g. 39], and update our beliefs about this function by evaluating $\mathbb{I}_{\phi\in\mathcal{S}}$ at a set of suitably chosen values for $\phi$. Whilst a GP does not enforce that $p(\phi)$ lies in $[0,1]$, our implementation below ensures that the mean of the GP does lie in this interval.
To implement this, we first need to choose a kernel function, $\boldsymbol{K}(\phi,\phi^{\prime})$, which specifies the covariance between each pair of $\phi$ values. We choose an exponential kernel $\boldsymbol{K}(\phi_{1},\phi_{2})=e^{-\frac{1}{2l^{2}}|\phi_{1}-\phi_{2}|}$, which is Markov. (This requires us to choose the value of $l$: we investigate this in Section C.3 of the Supplementary Material, where we find that at long as $l$ is at least 1, the value we use does not have a noticeable effect on the results.) If we have evaluated $\mathbb{I}_{\phi\in\mathcal{S}}$ at $\phi_{1}<\cdots<\phi_{N}$, and $\phi_{i}<\phi<\phi_{i+1}$ the Markov property means that our Gaussian approximation to $p(\phi)$ will only depend on the values of $\mathbb{I}_{\phi_{i}\in\mathcal{S}}$ and $\mathbb{I}_{\phi_{i+1}\in\mathcal{S}}$. Furthermore, for such a kernel the mean of the Gaussian process will be an average of these values and the prior mean – and thus will always lie within $[0,1]$.
@@
We can then calculate the estimate of $p(\phi)$ at discrete locations $\boldsymbol{\phi}^{\prime}=\{\phi_{1}^{\prime},\ldots,\phi_{N}^{\prime}\}$, as the mean of our GP:
$\hat{\boldsymbol{p}}(\boldsymbol{\phi}^{\prime})=\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi}^{\prime})^{\top}\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi})^{-1}\boldsymbol{p}(\boldsymbol{\phi}),$
where we use $\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi}^{\prime})$ to denote the matrix whose $(i,j)$th entry is $K(\phi_{i},\phi^{\prime}_{j})$. The most expensive step here computationally is calculating the inverse of the $N\times N$ matrix $\boldsymbol{K}(\boldsymbol{\phi},\boldsymbol{\phi})$, but the Markov property of our kernel means that this scales linearly with $N$.
Let $\pi(\phi)$ denote the unconditional distribution of $\phi$, i.e. $Beta(\frac{h}{2},\frac{h}{2})$. Let $q(\phi)$ denote the conditional distribution of $\phi|\phi\in\mathcal{S}$:
$\begin{split}q(\phi)&\propto\pi(\phi)\cdot\mathbb{I}_{\phi\in\mathcal{S}}\end{split}$
We can estimate this by replacing $\mathbb{I}_{\phi\in\mathcal{S}}$ with the mean of this quantity under our GP model. This is just the mean of the GP, and gives the estimate, $\hat{q}(\phi)$, where
@@
$p=\Pr(\phi\leq\phi_{lower}\text{ or }\phi\geq\phi_{upper}~{}|~{}\phi\in\mathcal{S})=q(\phi\leq\phi_{lower}\text{ or }\phi\geq\phi_{upper}).$ (10)
@@
Alternatively, we can use importance sampling with $\hat{q}$ as the proposal distribution. This means that we sample $\tilde{\phi}_{1},\ldots,\tilde{\phi}_{\tilde{N}}\sim\hat{q}(\phi)$; then
$\hat{p}=\frac{\sum_{j=1}^{\tilde{N}}\mathbb{I}\left(\tilde{\phi}_{j}\in\mathcal{S}~{}\text{and}~{}\left(\tilde{\phi}_{j}\leq\phi_{lower}\text{ or }\tilde{\phi}_{j}\geq\phi_{upper}\right)\right)\frac{\pi(\tilde{\phi}_{j})}{\hat{q}(\tilde{\phi}_{j})}}{\sum_{j=1}^{\tilde{N}}\mathbb{I}\left(\tilde{\phi}_{j}\in\mathcal{S}\right)\frac{\pi(\tilde{\phi}_{j})}{\hat{q}(\tilde{\phi}_{j})}}.$ (11)
Since we have already calculated $\mathbb{I}_{\phi\in\mathcal{S}}$ for the $\phi$ used to fit the GP, we can also include these points in our importance sampling estimator.
@@ 4.1 Different null hypotheses
There are alternative choices for the null hypothesis we may want to test, in particular to deal with the possibility of detected changepoints being close together. For example, we could decide to truncate the region if we find another changepoint within $h$ of the changepoint being tested. Alternatively, if we denote our changepoint of interest as $\hat{\tau}_{j}$, we can take the null hypothesis that there is no changepoint within $\{\hat{\tau}_{j-1}+1,\ldots,\hat{\tau}_{j+1}-1\}$ (by convention setting $\hat{\tau}_{0}=0$ and $\hat{\tau}_{K+1}=T$, where $K$ is the number of changepoints), or define the region we test to that closer to $\hat{\tau}_{j}$ than any other detected changepoint [see 22, 7].
@@ 4.2 Different changepoint detection methods
We have focused on two approaches to post-selection inference for the change in variance model. The CUSUM approach (Section 3.1) depends on being able to write down a set of linear inequalities in $\phi$ which can be resolved to obtain the set $\mathcal{S}$. We have focused on doing this for binary segmentation; however, it can be easily shown that we can do the same for variants – such as wild binary segmentation – which also involve maximizing $G_{s,e}(t)$ on a series of intervals. For our second approach (Section 3.2), we can use any method for changepoint estimation, since we just re-apply the algorithm for each sampled $\phi$.
For both methods, we will generate data sets for different values of $\phi$ and re-run our changepoint detection method on each data set. This is the main computational cost of both methods, though they differ in terms of the number of new data sets that have to be analysed. In the former method this is equal to the number of intervals of $\phi$ that correspond to a different solution path of e.g. binary segmentation. In most examples we have considered this is fewer than 10. For the second method the number of data sets depends on the Monte Carlo accuracy needed when estimating the $p$-value but tends to be of the order of 100s.
Often one can reduce the computational cost through early stopping when we run the changepoint detection algorithm. If we are conditioning on $\hat{\tau}\in\mathcal{M}$ then we can stop once this condition holds. By comparison if we are conditioning on $\mathcal{M}$ we can stop if ever we detect a change not in $\mathcal{M}$ for our new data. Also we can often re-use calculations. For example, to implement methods such as wild binary segmentation and narrowest over threshold we have to calculate the likelihood ratio statistic on a set of intervals $[s,e]\subseteq[1,T]$. The set of intervals is the same each time we implement the algorithm on $\boldsymbol{X}^{\prime}(\phi)$. Since $\Lambda_{s,e}(t,\phi)$ is constant on intervals which do not overlap with the region of interest (i.e., $e\leq\hat{\tau}-h$ or $s\geq\hat{\tau}+h+1$), for these intervals we only have to calculate $\Lambda_{s,e}(t)$ once. Therefore, at each sampling iteration we only have to re-calculate $\Lambda_{s,e}(t,\phi)$ on intervals which overlap with the region of interest.
@@ 4.3 Increasing power by conditioning on less information
where here both $\phi$ and $\boldsymbol{W}$ are allowed to vary.
@@ 5.1 Detecting changes using CUSUM
$\sigma_{2}^{2}=1$(a)$\sigma_{2}^{2}=0.25$(b)$\sigma_{2}^{2}=2$(c)$\sigma_{2}^{2}=4$(d)
Figure 7: QQ plots of $p$-values obtained using the CUSUM method. We set $T=200$, and sampled data under $H_{0}$ (no change) and with a single changepoint sampled randomly from $1,\ldots,T$. The variance before the change was $\sigma_{1}^{2}=1$ in each case; the variance after the change is $\sigma^{2}\in\{1,0.25,2,4\}$. In each case we estimate a single changepoint using binary segmentation. We calculated $p$-values for different values of $h$, including the case where the window size is equal to the whole data set.
Figure 7 shows QQ plots of $p$-values obtained by sampling from a model with $T=200$. In (a) we sample from $H_{0}$; in (b)–(d) we sample from a model with a single change at $\tau=100$, where in each case the variance before the changepoint is 1, and the variance after the changepoint is 0.25, 2, 4. We see that under $H_{0}$, $p\sim U(0,1)$, and under $H_{1}$ the test has power to detect changes, which increases with the window size $h$ and the size of the change.
@@
To see what happens when we have multiple changes, Figure 8 shows similar plots where we simulate from a model with 4 changes, with $T=200$ and $h=10,20$. In panels (a) and (b) we simulated from a model with 4 equally spaced changes, $\tau=40,80,120,160$. In (c) and (d), changepoints were chosen uniformly from $(h,T-h)$, with the constraint that the minimum distance between changepoints was at least $h$.
@@ 5.2 Detecting changes using the likelihood-ratio test
$N=50$$N=100$$N=200$$H_{0}$$\tau_{1}=\frac{T}{4},\tau_{2}=\frac{T}{2}$,$\tau_{3}=\frac{3T}{4}$$\tau_{1},\tau_{2},\tau_{3}$random
Figure 9: QQ plots of $p$-values obtained using the likelihood ratio for three scenarios – no change, three equally spaced changes, and three random changes – using binary segmentation to estimate changepoints. Each row of the figure corresponds to a scenario, and columns correspond to the number of samples used. Each line corresponds to the proportion of samples that were used to fit the Gaussian process.
Figure 9 shows QQ plots of estimated $p$-values we obtain on simulated data for three different scenarios: (1) no change; (2) three changes at fixed equidistant locations, $\tau=\frac{T}{4},\frac{T}{2},\frac{3T}{4}$; (3) three changes at random locations, sampled uniformly from $\{h,\ldots,T-h\}$ and constrained to be at least $h$ apart. In each case we set $T=200$ and $h=20$. For the cases with three changes, the variances on each segment were $(1,4,1,0.25)$. For each simulated data set, we estimated changepoints using binary segmentation (Figure 9) and calculated an estimate of the $p$-value corresponding to the first detected changepoint using a Gaussian process (as outlined in Section 3.2.2), and using a Gaussian process followed by importance sampling from the posterior distribution, for $N=50,100,200$.
@@ 6 Application to financial data
Figure 12: Application of our method to financial log returns data. Changepoints are estimated using (i) CUSUM statistic with binary segmentation (top); (ii) likelihood ratio statistic with binary segmentation (middle); (iii) likelihood ratio statistic with PELT [25] (bottom). In each case we estimated 11 changepoints. Vertical lines indicate changepoints: those with a $p$-value smaller than 0.05 (after controlling for multiple testing using the Holm-Bonferroni method) are shown in red, and those with a $p$-value larger than 0.05 in grey.
To investigate how well our method works in a real-world setting, we applied our method to financial data (Figure 12). The data is log returns from S&P500 stock market data for 2264 days from July 2010 to October 2019111The data was downloaded from https://www.kaggle.com/datasets/mathisjander/s-and-p500-volatility-prediction-time-series-data. We can see from Figure 12 that the data appears to have zero mean and several changes in variance. We estimated changepoints using three different methods: binary segmentation with the CUSUM statistic (top) and the likelihood ratio statistic (middle) and the penalized likelihood method PELT [25] (bottom), and calculated $p$-values for each estimated changepoint. We set thresholds so that in each case we estimated 11 changepoints, and set $h=50$. For the cases where we had to estimate $p$-values, we used a Gaussian process with $N=100$. On each plot estimated changepoints are shown by vertical lines, with red lines corresponding to estimated changepoint locations whose $p$-values were smaller than $0.05$ (after controlling for multiple testing using the Holm-Bonferroni method) and grey lines corresponding to $p$-values which were not significant. Using the likelihood ratio statistic generally seems to give better results than CUSUM, where there are more instances of multiple changepoints being estimated close together. However, these spurious changepoints generally have non-significant $p$-values associated with them, showing that our method for post-selection inference is effective at distinguishing true from false changepoint estimates.
@@ 7 Discussion
The strategies we have developed using Gaussian processes and importance sampling may be useful for other similar situations where we conduct post-selection inference but cannot characterise the selection event explicitly. More efficient implementations may be possible, for example using Bayesian optimisation [15] to adaptively choose $\phi$ values to consider based on some expected improvement in some appropriate measure of accuracy of the estimate of the $p$-value.
Acknowledgements
This work was supported by EPSRC grant number EP/V053590/1.